OpenStudy (anonymous):

How to simplify: √p^6 + √9p^6 + 3√p^6 - √p^2 Everything is inside of the radical

OpenStudy (anonymous):

\[\sqrt{p^6}= p^3 ; \sqrt{9p^6}=3p^3; 3\sqrt{p^6}=3p^3 ; \sqrt{p^2}=p; now solve..\]

OpenStudy (anonymous):

7p^3 - p ?

OpenStudy (anonymous):

yes u got it ...

OpenStudy (anonymous):

thank you!